OpenStudy (anonymous):

Amistre64 can you help me again?

OpenStudy (anonymous):

\[22xe ^{x ^{2}}/\sqrt{e ^{x ^{2}}}\]

OpenStudy (amistre64):

ok .... and this is a part of?

OpenStudy (anonymous):

I need to find what it equals. I know what the answer is again, but again do not know how they got it.

OpenStudy (anonymous):

want a screen shot?

OpenStudy (amistre64):

let me see the problem to get an better view of what you need; yes screen shot

OpenStudy (amistre64):

who is thinking up these things lol

OpenStudy (anonymous):

I told you.,,it's the same guy that thinks he's smarter than me! Telling me I'm wrong all the time!! Humph!

OpenStudy (amistre64):

\[\frac{22x.e^{x^2}}{\sqrt{e^{x^2 }}}\]

OpenStudy (amistre64):

and you wanna know how they get their answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just index laws

OpenStudy (amistre64):

lets clean it up some: \[\frac{22x.b}{\sqrt{b}}\] thats easier to see right? \[\frac{22x.b}{\sqrt{b}}*\frac{\sqrt{b}}{\sqrt{b}}=\frac{22x.b.b^{1/2}}{b}\] right?

OpenStudy (amistre64):

does \({b\over b}\) cancel to 1?

OpenStudy (amistre64):

leaves us: \[22x.b^{1/2}\] and we said \(b = e^{x^2}\) so rewrite it: \[22x.e^{(1/2)x^2}\]

OpenStudy (anonymous):

yes but l;et me try to catch up...had to go away to give some medicine to my son. Let me scroll up to look at what you've typed.

OpenStudy (anonymous):

You are the best...again thank you! You turned on my light bulb!

OpenStudy (amistre64):

lol...youre welcome